# Estimating Causal Effects for Binary Outcomes Using Per-Decision Inverse Probability Weighting

### v2

@@ Appendix F Proof of Equation (10) in Simulation
$=1-P(R_{t+1}(\bar{A}_{t-1},1,\bar{0})=0,R_{t+2}(\bar{A}_{t-1},1,\bar{0})=0,\ldots,R_{t+\Delta}(\bar{A}_{t-1},1,\bar{0})=0\mid H_{t})$ (47)
$=1-P(R_{t+1}(\bar{A}_{t-1},1,\bar{0})=0\mid H_{t})\cdot P(R_{t+2}(\bar{A}_{t-1},1,\bar{0})=0\mid H_{t})\cdots P(R_{t+\Delta}(\bar{A}_{t-1},1,\bar{0})=0\mid H_{t}),$ (48)
where (47) follows from Assumption 4, and (48) follows from the fact that $R_{t+1},\ldots,R_{t+\Delta}$ are independent of each other in the generative model. Similarly, for $E(Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0})$ we have
@@
$=1-P(R_{t+1}(\bar{A}_{t-1},0,\bar{0})=0\mid H_{t})\cdot P(R_{t+2}(\bar{A}_{t-1},0,\bar{0})=0\mid H_{t})\cdots P(R_{t+\Delta}(\bar{A}_{t-1},0,\bar{0})=0\mid H_{t}).$ (49)
@@
$=E\{E\left(\mathbbm{1}\left(R_{t+j+1}=0\right)\left|H_{t+j},\left.A_{t+j}=0\right)\right|H_{t},A_{t}=a\right\}$ (50)
@@
$=3/C\cdot\gamma^{1/\Delta},$ (52)
where (50) follows from Assumption 1 and the fact that $R_{t+j+1}$ only depends on $(Z_{t+j},A_{t+j})$ in the generative model. Plugging (52) into (48) and we have
@@
=$=\{1-\gamma^{(\Delta+0.5-0.5Z_{t})/\Delta}\cdot(3/C)^{\Delta-1}\}\cdot\exp\{0.1+0.2Z_{t}\}$ (53)
Plugging (52) into (49) and we have
@@
$=1-\gamma^{(\Delta+0.5-0.5Z_{t})/\Delta}\cdot(3/C)^{\Delta-1}$ (54)
Based on (53) and (54), we have:
@@
Appendix G Generalized Estimating Equations (GEE) Used in Simulation
@@
Because we are considering causal effects on the relative risk scale, the GEE assumes the following mean model using a log link: $E(Y_{t,\Delta}\mid H_{t},A_{t})=e^{g(H_{t})^{T}\alpha+A_{t}S_{t}^{T}\beta}=:\mu_{t}(\alpha,\beta)$. Because $Y_{t,\Delta}$ is a binary variable, its variance is associated with the mean model: $\text{var}(Y_{t,\Delta}\mid H_{t},A_{t})=\mu_{t}(\alpha,\beta)\{1-\mu_{t}(\alpha,\beta)\}$. The estimating equation can be constructed following the original GEE proposal in [30]. Below we write down the estimating equation for GEE.ind:
@@
Appendix H Generalization to Other Reference Regimes
@@
As pointed out by a reviewer, the choice of the reference treatment regime depends on the scientific question, and in some applications other reference regimes may be of interest. For example, in [18] they considered causal excursion effect contrasting treatment trajectories $(\bar{A}_{t-1},1,\bar{A}_{t+1:t+\Delta-1})$ and $(\bar{A}_{t-1},1,\bar{A}_{t+1:t+\Delta-1})$, where $\bar{A}_{t+1:t+\Delta-1}:=(A_{t+1},A_{t+2},\ldots,A_{t+\Delta-1})$. That is, they considered excursions starting from decision point $t$ that receive treatment (or no treatment) at $t$, and then switch back to follow the MRT policy for the next $\Delta-1$ decision points.
We generalize our method to reference regimes that are in between the reference regime we considered in the main paper and the reference regime considered in [18]. That is, excursions starting from decision point $t$ that receive treatment (or no treatment) at $t$, then receive no treatment for the next $K$ decision points (with some $10\leq K\leq\Delta-$ 1), then follow the MRT policy for the next $\Delta-1-K$ decision points.
@@
$\text{CEE}^{(K)}\left\{S_{t}(\bar{A}_{t-1})\right\}=\log\frac{E\left\{Y_{t,\Delta}(\bar{A}_{t-1},1,\bar{0}_{K},\bar{A}_{t+K+1:t+\Delta-1})\mid S_{t}(\bar{A}_{t-1}),I_{t}(\bar{A}_{t-1})=1\right\}}{E\left\{Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{K},\bar{A}_{t+K+1:t+\Delta-1})\mid S_{t}(\bar{A}_{t-1}),I_{t}(\bar{A}_{t-1})=1\right\}}.$ (55)
@@
$\text{CEE}^{(K)}\left\{S_{t}(\bar{A}_{t-1})\right\}=\log\frac{E\bigg{[}E\left\{\prod_{j=t+1}^{t+K}\left[\frac{\mathbbm{1}(A_{j}=0)}{1-p_{j}\left(H_{j}\right)}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=1,H_{t},I_{t}=1\right\}\mid S_{t},I_{t}=1\bigg{]}}{E\bigg{[}E\left\{\prod_{j=t+1}^{t+K}\left[\frac{\mathbbm{1}(A_{j}=0)}{1-p_{j}\left(H_{j}\right)}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=0,H_{t},I_{t}=1\right\}\mid S_{t},I_{t}=1\bigg{]}}.$ (56)
The difference between the two identification results (56) and (4) is that in (56) the product of weights is only up to decision point $t+K$ instead of up to $t+\Delta-1$, because after decision point $t+K$ the treatment trajectory comes back to the MRT protocol and thus no additional weight is needed.
@@
$U_{i}^{(K)}(\alpha,\beta^{(K)})=\sum_{t=1}^{T}I_{it}e^{-A_{it}S_{it}^{T}\beta^{(K)}}\epsilon_{it}(\alpha,\beta^{(K)})M_{it}W_{it}^{(K)}\begin{bmatrix}g(H_{it})\\ \{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}\end{bmatrix},$ (57)
@@
Below we present simulation results to compare this generalized version of our method (generalized pd-EMEE) and a generalized version of EMEE by [11], and see how the relative efficiency depends on the choice of reference regime. Here the EMEE by [11] needs to be generalized to handle the reference regimes we consider here. In particular, for the generalized version of EMEE, the estimating function would be the same as 57 except that $W_{it}^{(K)}$ would be replaced by

### v3

@@ H.1 Generative Model used in Simulation
$=1-P(R_{t+1}(\bar{A}_{t-1},1,\bar{0})=0,R_{t+2}(\bar{A}_{t-1},1,\bar{0})=0,\ldots,R_{t+\Delta}(\bar{A}_{t-1},1,\bar{0})=0\mid H_{t})$ (H.5)
$=1-P(R_{t+1}(\bar{A}_{t-1},1,\bar{0})=0\mid H_{t})\cdot P(R_{t+2}(\bar{A}_{t-1},1,\bar{0})=0\mid H_{t})\cdots P(R_{t+\Delta}(\bar{A}_{t-1},1,\bar{0})=0\mid H_{t}),$ (H.6)
where (H.5) follows from Assumption 4, and (H.6) follows from the fact that $R_{t+1},\ldots,R_{t+\Delta}$ are independent of each other in the generative model. Similarly, for $E(Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0})$ we have
@@
$=1-P(R_{t+1}(\bar{A}_{t-1},0,\bar{0})=0\mid H_{t})\cdot P(R_{t+2}(\bar{A}_{t-1},0,\bar{0})=0\mid H_{t})\cdots P(R_{t+\Delta}(\bar{A}_{t-1},0,\bar{0})=0\mid H_{t}).$ (H.7)
@@
$=E\{E\left(\mathbbm{1}\left(R_{t+j+1}=0\right)\left|H_{t+j},\left.A_{t+j}=0\right)\right|H_{t},A_{t}=a\right\}$ (H.8)
@@
$=3/C\cdot\gamma^{1/\Delta},$ (H.10)
where (H.8) follows from Assumption 1 and the fact that $R_{t+j+1}$ only depends on $(Z_{t+j},A_{t+j})$ in the generative model. Plugging (H.10) into (H.6) and we have
@@
=$=\{1-\gamma^{(\Delta+0.5-0.5Z_{t})/\Delta}\cdot(3/C)^{\Delta-1}\}\cdot\exp\{0.1+0.2Z_{t}\}$ (H.11)
Plugging (H.10) into (H.7) and we have
@@
$=1-\gamma^{(\Delta+0.5-0.5Z_{t})/\Delta}\cdot(3/C)^{\Delta-1}$ (H.12)
Based on (H.11) and (H.12), we have:
@@
Appendix I Generalized Estimating Equations (GEE) Used in Simulation
@@
Because we are considering causal effects on the relative risk scale, the GEE assumes the following mean model using a log link: $E(Y_{t,\Delta}\mid H_{t},A_{t})=e^{g(H_{t})^{T}\alpha+A_{t}S_{t}^{T}\beta}=:\mu_{t}(\alpha,\beta)$. Because $Y_{t,\Delta}$ is a binary variable, its variance is associated with the mean model: $\text{var}(Y_{t,\Delta}\mid H_{t},A_{t})=\mu_{t}(\alpha,\beta)\{1-\mu_{t}(\alpha,\beta)\}$. The estimating equation can be constructed following the original GEE proposal in Liang and Zeger (1986). Below we write down the estimating equation for GEE.ind:
@@
Appendix J Generalization to Other Reference Regimes
@@
As pointed out by a reviewer, the choice of the reference treatment regime depends on the scientific question, and in some applications other reference regimes may be of interest. For example, in Boruvka and others (2018) they considered causal excursion effect contrasting treatment trajectories $(\bar{A}_{t-1},1,\bar{A}_{t+1:t+\Delta-1})$ and $(\bar{A}_{t-1},1,\bar{A}_{t+1:t+\Delta-1})$, where $\bar{A}_{t+1:t+\Delta-1}:=(A_{t+1},A_{t+2},\ldots,A_{t+\Delta-1})$. That is, they considered excursions starting from decision point $t$ that receive treatment (or no treatment) at $t$, and then switch back to follow the MRT policy for the next $\Delta-1$ decision points.
We generalize our method to reference regimes that are in between the reference regime we considered in the main paper and the reference regime considered in Boruvka and others (2018). That is, excursions starting from decision point $t$ that receive treatment (or no treatment) at $t$, then receive no treatment for the next $K$ decision points (with some $10\leq K\leq\Delta-$ 1), then follow the MRT policy for the next $\Delta-1-K$ decision points.
@@
$\text{CEE}^{(K)}\left\{S_{t}(\bar{A}_{t-1})\right\}=\log\frac{E\left\{Y_{t,\Delta}(\bar{A}_{t-1},1,\bar{0}_{K},\bar{A}_{t+K+1:t+\Delta-1})\mid S_{t}(\bar{A}_{t-1}),I_{t}(\bar{A}_{t-1})=1\right\}}{E\left\{Y_{t,\Delta}(\bar{A}_{t-1},0,\bar{0}_{K},\bar{A}_{t+K+1:t+\Delta-1})\mid S_{t}(\bar{A}_{t-1}),I_{t}(\bar{A}_{t-1})=1\right\}}.$ (J.1)
@@
$\text{CEE}^{(K)}\left\{S_{t}(\bar{A}_{t-1})\right\}=\log\frac{E\bigg{[}E\left\{\prod_{j=t+1}^{t+K}\left[\frac{\mathbbm{1}(A_{j}=0)}{1-p_{j}\left(H_{j}\right)}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=1,H_{t},I_{t}=1\right\}\mid S_{t},I_{t}=1\bigg{]}}{E\bigg{[}E\left\{\prod_{j=t+1}^{t+K}\left[\frac{\mathbbm{1}(A_{j}=0)}{1-p_{j}\left(H_{j}\right)}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=0,H_{t},I_{t}=1\right\}\mid S_{t},I_{t}=1\bigg{]}}.$ (J.2)
The difference between the two identification results (J.2) and (4) is that in (J.2) the product of weights is only up to decision point $t+K$ instead of up to $t+\Delta-1$, because after decision point $t+K$ the treatment trajectory comes back to the MRT protocol and thus no additional weight is needed.
@@
$U_{i}^{(K)}(\alpha,\beta^{(K)})=\sum_{t=1}^{T}I_{it}e^{-A_{it}S_{it}^{T}\beta^{(K)}}\epsilon_{it}(\alpha,\beta^{(K)})M_{it}W_{it}^{(K)}\begin{bmatrix}g(H_{it})\\ \{A_{it}-\tilde{p}_{t}(S_{it})\}S_{it}\end{bmatrix},$ (J.3)
@@
Below we present simulation results to compare this generalized version of our method (generalized pd-EMEE) and a generalized version of EMEE by Qian and others (2021), and see how the relative efficiency depends on the choice of reference regime. Here the EMEE by Qian and others (2021) needs to be generalized to handle the reference regimes we consider here. In particular, for the generalized version of EMEE, the estimating function would be the same as J.3 except that $W_{it}^{(K)}$ would be replaced by